# Variational Learning of Disentangled Representations

### v1

@@ 4 Comparison to previous approaches
VAEs (Kingma and Welling, 2014) are generative models that learn latent representations by maximizing the evidence lower bound (ELBO) on the data log-likelihood:
$\mathbb{E}_{q_{z|x}}[\log p(x\mid z)]-\operatorname{D_{\mathrm{KL}}}\left(q_{z\mid x}\ \|\ p_{z}\right)\leq\log p(x),$
@@
Conditional VAEs   (Sohn et al., 2015) incorporate supervision to the standard VAE model by conditioning both the encoder and decoder on an observed label $y$, yielding the following objective:
$\mathbb{E}_{q_{z|x,y}}[\log p(x\mid z,y)]-\operatorname{D_{\mathrm{KL}}}\left(q(z\mid x,y)\|p(z)\right)\leq\log p(x|y).$
@@
Conditional Subspace VAEs  (CSVAEs) (Klys et al., 2018), explicitly factorize the latent space into a shared component $z$ a label-specific component $w$ (see Supplementary Figure 1(a)). Similarly, their hierarchical extension Beker et al. (2024) introduces an intermediate latent variable between $x$ and $(z,w)$. As in our approach, to encourage disentanglement, CSVAEs introduce an adversarial regularization term that penalizes mutual information between $z$ and $y$, thereby discouraging predictability of $y$ from $z$. They are learned by optimizing the following lower bound on $\log p(x\mid y)$:
[…]
Domain Invariant VAEs (DIVA) and Characteristic-capturing VAEs (CCVAE)   DIVA (Ilse et al., 2020) and CCVAE (Joy et al., 2020), shown in Supplementary Figure 1(b), introduce two latent variables, $z$ and $w$, where $w$ captures label-related features by jointly optimizing a classifier $p(w\mid y)$ jointly alongside the remaining objective. For fully supervised cases, the DIVA model optimizes
$\textstyle\mathbb{E}_{q_{z,w|x}}[\log p(x\mid z,w)]+\mathbb{E}_{q_{w|x}}[\log q(y\mid w)]-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}\|p_{z}\right)-\operatorname{D_{\mathrm{KL}}}\left(q_{w\mid x}\|p_{w\mid y}\right)\leq\log p(x,y).$
@@
$\textstyle\mathbb{E}_{q_{z,w|x,y}}\left[\frac{q(y\mid w)}{q(y\mid x)}\log\frac{p(x\mid z,w)}{q(y\mid w)}\right]-\operatorname{D_{\mathrm{KL}}}\left(q_{z|x}\|p_{z|y}\right)-\operatorname{D_{\mathrm{KL}}}\left(q_{w\mid x}\|p_{w\mid y}\right)+\log q(y\mid x).$
Comparison:   In contrast to our approach, which relies on minimal independence Assumptions (1 -2), CSVAEs, DIVA, and CCVAE incorporate stronger structural assumptions directly into their objectives. Additionally, DIVA and CCVAE lack explicit mechanisms to remove residual dependence between $z$ and $y$, often leading to information leakage into the shared representation.
@@ 5 Experiments
We evaluate DISCoVeR with existing methods on both synthetic datasets and real datasets. When applicable, we assess reconstruction quality using negative log-likelihood (NLL), root mean squared error (RMSE), and the absolute deviation from the optimal Bayes classifier on the completed reconstruction, denoted as $\Delta$ – Bayes. We measure disentanglement using a neural estimation Belghazi et al. (2018) of the the mutual information $I(z;w)$. For all experiments, model architectures, hyper-parameters and all other implementation details are provided in Appendix E.
@@ 5.1.1 Parametric model
Results: Table 1 shows that DISCoVeR (ours) best approximates the analytic posteriors, resulting in the lowest deviation from the optimal Bayes classifier and the best reconstruction.
Table 1: Parametric model results: DISCoVeR (ours) outperforms all competitors across all metrics.
NLL $\downarrow$ $\operatorname{D_{\mathrm{KL}}}(q_{z\mid x}\;||\>p_{z\mid x})\downarrow$ $\operatorname{D_{\mathrm{KL}}}(q_{w\mid x}\;||\>p_{w\mid x})\downarrow$ $\Delta$ – Bayes $\downarrow$
@@
DISCoVeR (ours) $\mathbf{1.769\pm 0.003}$ $\mathbf{0.17\pm 0.01}$ $\mathbf{10.10\pm 0.73}$ $\mathbf{0.1\pm 0.28}$
@@ 5.1.2 Noisy Swiss roll
Dataset:   We use a noisy variant of the labeled Swiss Roll dataset (Marsland, 2014; Klys et al., 2018), generating $n=20,000$ and assigning binary labels based on a lengthwise split, with labels flipped at rate $\rho$. The common geometry (its projection along the 2D plane) remains intact, he conditional structure along the third axis becomes noisy. Figure 2A illustrates the setup.
Table 2: Noisy Swiss roll results: DISCoVeR  yields lowest deviation from optimal-Bayes, while maintaining low latent leakage and high reconstruction accuracy.
@@
CSVAE - N.A. $0.011$ $3.303\pm 0.003$ $23.88\pm 12.02$
CSVAE $\mathbf{0.002}$ $\mathbf{3.302\pm 0.003}$ $17.99\pm 14.58$
HCSVAE - N.A. $0.062$ $\mathbf{3.302\pm 0.002}$ $30.00\pm 0.00$
HCSVAE $0.003$ $\mathbf{3.302\pm 0.002}$ $30.00\pm 0.00$
DIVA $0.423$ $\mathbf{3.302\pm 0.003}$ $1.88\pm 1.05$
CCVAE $0.254$ $\mathbf{3.302\pm 0.002}$ $2.21\pm 0.84$
DISCoVeR (Ours) $0.004$ $\mathbf{3.302\pm 0.002}$ $\mathbf{1.14\pm 0.21}$
Optimal disentanglement:   Since the Swiss Roll is sliced at the center and label noise is applied uniformly, marginalizing over labels yields a symmetric spiral centered along the roll—i.e., the marginal posterior $p(z\mid x)$ is label-invariant. In contrast, the conditional component retains a noisy but informative signal, with a uniform noise level of $\rho=0.3$. As a result, the Bayes optimal classifier trained on any realistic representation is upper-bounded at 70% accuracy.
Results:   Figure 2 presents qualitative and quantitative results, showing that DISCoVeR both models the label noise accurately and effectively disentangles shared and condition-specific structure. Notably, DISCoVeR captures the marginal data distribution, successfully recovering the expected spiral pattern, as shown in Figure 2B (right). Additionally, the results in Table 2 show that DISCoVeR achieves the lowest deviation from the optimal Bayes classifier and minimal information leakage between latent variables, while preserving reconstruction quality.
Figure 2: A: Noisy labeled Swiss Roll dataset. B (left): Reconstruction and conditional embedding from DISCoVeR. B (right): Marginal reconstruction from the shared embedding, recovering the spiral structure. C: Comparison of reconstruction and conditional embeddings across models.
@@ 5.2.1 Noisy colored MNIST
Dataset:   We use a modified MNIST (Deng, 2012) dataset constructed from $60,000$ duplicated images: in one copy we remove the red channel ($y=0$) and in the other we remove the green channel ($y=1$). The digit shape remains intact following the blue channel. Label noise is introduced by flipping labels with probability $\rho\in\{0,0.1,0.2,0.3,0.4\}$.
@@
Results:   We evaluate marginal coloring reconstruction by DISCoVeR and previous methods. Under no label noise ($\rho=0$), all methods perform similarly (Supplementary Figure 3). At all non-zero noise levels, DISCoVeR consistently outperforms competitors, whose marginal reconstructions are averaged over class-conditioned outputs. Metrics for $\rho=0.3$ are shown in Supplementary Table 2, with results for other noise levels in Supplementary Figure 4.
Figure 3: Colored MNIST results for noise level $\rho=0.3$. DISCoVeR is the only model that consistently reconstructs the correct constant color matching the true marginal $p(x|B)$.
@@ 5.2.2 CelebA-Glasses
Dataset:   We use all CelebA (Liu et al., 2015) images labeled with eyeglasses attribute ($y=1$), and twice as many randomly sampled images without ($y=0$), totaling $n=35,712$ samples.
Results:   Figure 4 shows that DISCoVeR accurately reconstructs input images while producing shared embeddings that marginalize over the eyeglasses attribute, consistently adding "pseudo-glasses" to all samples. Competing methods are shown in Supplementary Figure 5, with quantitative results in Supplementary Table 3. While reconstruction quality is comparable across methods, DISCoVeR achieves significantly better disentanglement.
Figure 4: CelebA-Eyeglasses results. Top: Original images with and without eyeglasses. Middle: Full reconstructions by DISCoVeR. Bottom: Reconstructions solely from common embeddings $z$.
@@ 5.2.3 Single cell RNA-sequencing (scRNA-seq) data from Lupus dataset
Figure 5: A, B (left): UMAP embeddings of gene counts from the IFN-$\beta$ stimulation dataset. A, B (middle): Common embedding $z$ preserves cell type-specific differences, while correcting for the stimulation effect. A, B (right): Conditional embedding $w$ captures the stimulation effect. C: Comparison of original and reconstructed log-mean (top) and log-variance (bottom) across cells for the two different reconstructions offered by DISCoVeR.
Results:   Supplementary Table 5 shows that DISCoVeR effectively achieves the desired behavior with strong empirical performance, where only cell type information is captured in $z$ (Figure 5A, middle) while the effects of IFN-$\beta$ stimulation are wholly represented in $w$ (Figure 5B, right). Other approaches either (1) achieve mixing in the $z$ space, but compromise on keeping cell types separated or (2) leak information about stimulation into the $z$ space (Supplementary Figure 6).
Facilitating interpretability:   By enabling marginalized reconstructions, DISCoVeR provides a direct link between shared embeddings and gene expression, offering clearer insight into the effects of IFN-$\beta$ stimulation, unlike other methods. In Figure 5C, comparing variance across marginal and full reconstructions accurately recovers gene-level differences associated with IFN-$\beta$ stimulation, including ISG15, FTL, CCL8, CXCL10, CXCL11, APOBEC3A, IL1RN, IFITM3 and RSAD2.
@@ 6 Conclusion
In this work we introduced a variational framework for disentangled representation learning that explicitly separates condition-invariant and condition-aware factors, while avoiding overly restrictive independence assumptions. Unlike prior work, DISCoVeR incorporates two reconstruction paths—one based solely on the shared latent variable $z$, and the other on the full pair $(z,w)$. Our model simultaneously learns informative shared representations, and captures structured variation across conditions. Experimental results demonstrate that DISCoVeR achieves strong reconstruction, minimal information leakage, and accurate modeling of conditional effects—consistently outperforming existing methods.
Our model assumes that the condition labels are fully observed and discrete. We leave extensions of our framework to semi-supervised or continuous condition settings for future work.
@@ A.1 Domain Generalization
The task of representation disentanglement is closely related to the field of domain generalization [Muandet et al., 2013], which assumes limited or no access to target domain samples and aims to learn representations that can be readily adapted, often via transfer learning, to new, unseen domains.
As noted by Wang et al. [2019], existing methods in domain generalization can be broadly categorized into two main approaches: (i) approaches for reducing the inter-domain differences, often by using adversarial techniques [Ghifary et al., 2015, Wang et al., 2017, Motiian et al., 2017, Li et al., 2018, Carlucci et al., 2019, Li et al., 2018, Wang et al., 2019, Akuzawa et al., 2020, Zhu et al., 2022, Gokhale et al., 2023, Dayal et al., 2023, Cheng et al., 2023, Chen et al., 2024], and (ii) Approaches that construct an ensemble of domain-specific models, and then fuse their representations to form a unified, domain-agnostic representation [Ding and Fu, 2017, Mancini et al., 2018, Zhou et al., 2021, Muhammad et al., 2024].
Additional strategies for domain generalization include contrastive learning approaches [Kim et al., 2021], methods based on distribution alignment via metrics [Muandet et al., 2013, Sun and Saenko, 2016], and techniques utilizing custom network architectures, for instance by incorporating domain-specific adapters between shared layers [Rebuffi et al., 2017, 2018, Li and Vasconcelos, 2019, Omi et al., 2022].

### v2

@@ 4 Comparison to previous approaches
VAEs (vae) are generative models that learn latent representations by maximizing the evidence lower bound (ELBO) on the data log-likelihood:
$\mathbb{E}_{q_{z|x}}[\log p(x\mid z)]-D_{\mathrm{KL}}\left(q_{z\mid x}\ \|\ p_{z}\right)\leq\log p(x),$
@@
Conditional VAEs   (CVAE) incorporate supervision to the standard VAE model by conditioning both the encoder and decoder on an observed label $y$, yielding the following objective:
$\mathbb{E}_{q_{z|x,y}}[\log p(x\mid z,y)]-D_{\mathrm{KL}}\left(q(z\mid x,y)\|p(z)\right)\leq\log p(x|y).$
@@
Conditional Subspace VAEs  (CSVAEs) (CSVAE), explicitly factorize the latent space into a shared component $z$ and a label-specific component $w$ (see Supplementary Figure 1(a)). Similarly, their hierarchical extension beker2024patches introduces an intermediate latent variable between $x$ and $(z,w)$. As in our approach, to encourage disentanglement, CSVAEs introduce an adversarial regularization term that penalizes mutual information between $z$ and $y$, thereby discouraging predictability of $y$ from $z$. They are learned by optimizing the following lower bound on $\log p(x\mid y)$:
$\textstyle\mathbb{E}_{q_{z,w|x,y}}[\log p(x\mid w,y)]\!-\!\mathbb{E}_{q_{z|x}}[\int g(y\mid z)\log g(y\mid z)\,dy]\!-\!D_{\mathrm{KL}}\left(q_{w\mid x,y}\,\|\,q_{w\mid y}\right)\!-\!D_{\mathrm{KL}}\left(q_{z\mid x}\,\|\,p_{z}\right).$
Domain Invariant VAEs (DIVA) and Characteristic-capturing VAEs (CCVAE)   DIVA (diva) and CCVAE (CCVAE), shown in Supplementary Figure 1(b), introduce two latent variables, $z$ and $w$, where $w$ captures label-related features by jointly optimizing a classifier $p(w\mid y)$ jointly alongside the remaining objective. For fully supervised cases, the DIVA model optimizes
$\textstyle\mathbb{E}_{q_{z,w|x}}[\log p(x\mid z,w)]+\mathbb{E}_{q_{w|x}}[\log q(y\mid w)]-D_{\mathrm{KL}}\left(q_{z|x}\|p_{z}\right)-D_{\mathrm{KL}}\left(q_{w\mid x}\|p_{w\mid y}\right)\leq\log p(x,y).$
@@
$\textstyle\mathbb{E}_{q_{z,w|x,y}}\left[\frac{q(y\mid w)}{q(y\mid x)}\log\frac{p(x\mid z,w)}{q(y\mid w)}\right]-D_{\mathrm{KL}}\left(q_{z|x}\|p_{z|y}\right)-D_{\mathrm{KL}}\left(q_{w\mid x}\|p_{w\mid y}\right)+\log q(y\mid x).$
Comparison:   In prior methods, reconstruction is performed jointly from both representations $z$ and $w$, via $p(x\mid z,w)$. This design provides no incentive to distribute information meaningfully between $z$ and $w$: the model can place all relevant information into $w$, leaving $z$ either uninformative or entangled with $w$. Our method addresses this limitation through two key components: (i) a separate reconstruction term from $z$ alone, which explicitly forces $z$ to capture informative, condition-invariant structure. Theoretically, we show that the additional term $p(\hat{x}\mid z)$ is required to bound the gap between our approximate posterior and the full model ELBO under the factorization assumption; (ii) a prior over $w$ conditioned on the mean of $z$, which, through label-specific aggregates, discourages leakage of class-invariant information into $w$ and reduces redundancy between the two representations.
Another novelty of our approach is a principled probabilistic objective that enforces the correct conditional independence. Without an explicit probabilistic model, prior methods sought to enforce $z\perp w\mid x,y$. In contrast, our formulation shows that the proper requirement is the weaker condition $z\perp w\mid y$. As established in Section 2.1, our model satisfies $z\perp w\mid y$, but not $z\perp w\mid x,y$.
To enforce this criterion, we propose a theoretically grounded variational objective that uses a conditional ELBO for the dependent representation $w$, with the prior over $w$ conditioned on the mean of the independent representation $z$.
@@ 5 Experiments
Datasets: We evaluate DisCoVR against existing approaches on synthetic data, natural images, and biological data. These datasets were chosen to probe condition-invariant structure and to ensure comparability with prior work: for instance, Swiss rolls and CelebA were used in CSVAE, and CelebA also in CCVAE.
Evaluation: When applicable, we evaluate reconstruction quality using negative log-likelihood (NLL), root mean squared error (RMSE), and the absolute deviation from the optimal-Bayes classifier on the reconstructed data, denoted as $\Delta$–Bayes. Disentanglement is quantified via a neural estimator of the mutual information $I(z;w)$ (pmlr-v80-belghazi18a). Full model architectures, hyperparameters, and additional implementation details are provided in Appendix H. Our results show that DisCoVR achieves superior performance across all experiments.
@@ 5.1.1 Parametric model
Results: Table 1 shows that DisCoVR (ours) best approximates the analytic posteriors, resulting in the lowest deviation from the optimal Bayes classifier and the best reconstruction.
Table 1: Parametric model results: DisCoVR (ours) outperforms all competitors across all metrics.
NLL $\downarrow$ $D_{\mathrm{KL}}(q_{z\mid x}\;||\>p_{z\mid x})\downarrow$ $D_{\mathrm{KL}}(q_{w\mid x,y}\;||\>p_{w\mid x,y})\downarrow$ $\Delta$ – Bayes $\downarrow$
@@
DisCoVR (ours) $\mathbf{1.769\pm 0.003}$ $\mathbf{0.17\pm 0.01}$ $\mathbf{10.10\pm 0.73}$ $\mathbf{0.1\pm 0.28}$
@@ 5.1.2 Noisy Swiss roll
Dataset:   We use a noisy variant of the labeled Swiss Roll dataset (swissroll; CSVAE), generating $n=20,000$ and assigning binary labels based on a lengthwise split, with labels flipped at rate $\rho$. The common geometry (its projection along the 2D plane) remains intact, the conditional structure along the third axis becomes noisy. Figure 2A illustrates the setup.
Table 2: Noisy Swiss roll results: DisCoVR (ours) yields lowest deviation from optimal-Bayes, maintains low latent leakage, and high reconstruction accuracy.
@@
CSVAE - N.A. $0.047\pm 0.023$ $3.303\pm 0.003$ $23.88\pm 12.02$
CSVAE $0.031\pm 0.025$ $\mathbf{3.302\pm 0.003}$ $17.99\pm 14.58$
HCSVAE - N.A. $0.024\pm 0.012$ $\mathbf{3.302\pm 0.002}$ $30.00\pm 0.00$
HCSVAE $\mathbf{0.002\pm 0.001}$ $\mathbf{3.302\pm 0.002}$ $30.00\pm 0.00$
DIVA $0.336\pm 0.083$ $\mathbf{3.302\pm 0.003}$ $1.88\pm 1.05$
CCVAE $0.502\pm 0.089$ $\mathbf{3.302\pm 0.002}$ $2.21\pm 0.84$
DisCoVR (Ours) $0.005\pm 0.002$ $\mathbf{3.302\pm 0.002}$ $\mathbf{1.14\pm 0.21}$
Optimal disentanglement:   Since the Swiss Roll is sliced at the center and label noise is applied uniformly, marginalizing over labels yields a symmetric spiral centered along the roll—i.e., the marginal posterior $p(z\mid x)$} is label-invariant. In contrast, the conditional component retains a noisy but informative signal, with a uniform noise level of $\rho=0.3$. As a result, the Bayes optimal classifier trained on any realistic representation is upper-bounded at 70% accuracy. Results:   Figure 2 presents qualitative and quantitative results, showing that DisCoVR both models the label noise accurately and effectively disentangles shared and condition-specific structure. Notably, DisCoVR captures the marginal data distribution, successfully recovering the expected spiral pattern, as shown in Figure 2B (right). Additionally, the results in Table 2 show that DisCoVR achieves the lowest deviation from the optimal Bayes classifier and minimal information leakage between latent variables, while preserving reconstruction quality.
Figure 2: A: Noisy labeled Swiss Roll dataset. B (left): Reconstruction and conditional embedding from DisCoVR. B (right): Marginal reconstruction from the shared embedding, recovering the spiral structure. C: Comparison of reconstruction and conditional embeddings across models.
@@ 5.2.1 Noisy colored MNIST
Dataset:   We use a modified MNIST (deng2012mnist) dataset constructed from $60,000$ duplicated images: in one copy we remove the red channel ($y=0$) and in the other we remove the green channel ($y=1$). The digit shape remains intact following the blue channel. Label noise is introduced by flipping labels with probability $\rho\in\{0,0.1,0.2,0.3,0.4\}$.
@@
Results:   We evaluate marginal coloring reconstruction by DisCoVR and previous methods. Under no label noise ($\rho=0$), all methods perform similarly (Supplementary Figure 3). At all non-zero noise levels, DisCoVR consistently outperforms competitors, whose marginal reconstructions are averaged over class-conditioned outputs. Metrics for $\rho=0.3$ are shown in Supplementary Table 2, with results for other noise levels in Supplementary Figure 4.
Figure 3: Colored MNIST reconstructions from the label-agnostic representation $z$ for noise level $\rho=0.3$. DisCoVR is the only model that consistently reconstructs mixed “semi-red/blue” tone (purple) indicating that color information has been removed and matching the true marginal $p(x|B)$.
@@ 5.2.2 CelebA-Glasses
Dataset:   We use all CelebA (liu2015faceattributes) images labeled with eyeglasses attribute ($y=1$), and twice as many randomly sampled images without ($y=0$), totaling $n=35,712$ samples.
Results:   Figure 4 shows that DisCoVR accurately reconstructs input images while producing shared embeddings that marginalize over the eyeglasses attribute, consistently adding ”pseudo-glasses” to all samples. Competing methods are shown in Supplementary Figure 5, with quantitative results in Supplementary Table 3. While reconstruction quality is comparable across methods, DisCoVR achieves significantly better disentanglement.
Figure 4: CelebA-Eyeglasses results. Top: Original images with/without eyeglasses. Middle: Full reconstructions by DisCoVR. Bottom: Reconstructions solely from common embeddings $z$. A shared representation needs to be invariant to $y$ (presence or absence of glasses). Indeed, all reconstructed faces display an intermediate ”semi-glasses” appearance, regardless of the original label.
Results for an additional experiment with the wearing-hat attribute are provided in Appendix G.
@@ 5.2.3 Single cell RNA-sequencing (scRNA-seq) data from Lupus dataset
Figure 5: A–B (left): UMAPs of raw gene counts from the IFN-$\beta$ dataset. A–B (middle): Shared embedding $z$ aligns cells by type while removing stimulation effects. A–B (right): Condition-specific embedding $w$ isolates the stimulation effect. C: Reconstructions from both $z$ and $w$ (yellow) recover empirical gene means and variances, while reconstructions from $z$ alone (blue) miss the stimulation-induced variance, confirming that $z$ discards $y$ while preserving cell-type features.
Results:   Supplementary Table 17 shows that DisCoVR effectively achieves the desired behavior with strong empirical performance, where only cell type information is captured in $z$ (Figure 5A, middle) while the effects of IFN-$\beta$ stimulation are wholly represented in $w$ (Figure 5B, right). Other approaches either (1) achieve mixing in the $z$ space, but compromise on keeping cell types separated or (2) leak information about stimulation into the $z$ space (Supplementary Figure 6).
Facilitating interpretability:   By enabling marginalized reconstructions, DisCoVR provides a direct link between shared embeddings and gene expression, offering clearer insight into the effects of IFN-$\beta$ stimulation, unlike other methods. In Figure 5C, comparing variance across marginal and full reconstructions accurately recovers gene-level differences associated with IFN-$\beta$ stimulation, including ISG15, FTL, CCL8, CXCL10, CXCL11, APOBEC3A, IL1RN, IFITM3 and RSAD2.
@@ 6 Conclusion
In this work we introduced a variational framework for disentangled representation learning. Our formulation explicitly separates condition-invariant and condition-aware factors. Unlike prior work, DisCoVR incorporates two reconstruction paths, one based solely on the shared latent variable $z$, and the other on both latent variables $z$ and $w$. Our model simultaneously learns informative shared representations, and captures structured variation across conditions. Experimental results demonstrate that DisCoVR achieves strong reconstruction, minimal information leakage, and accurate modeling of conditional effects, consistently outperforming existing methods.
@@ A.1 Domain generalization
The task of representation disentanglement is closely related to the field of domain generalization (muandet2013domain), which assumes limited or no access to target domain samples and aims to learn representations that can be readily adapted, often via transfer learning, to new, unseen domains.
As noted by wang2019learning, existing methods in domain generalization can be broadly categorized into two main approaches: (i) approaches for reducing the inter-domain differences, often by using adversarial techniques (ghifary2015domain; wang2017select; motiian2017unified; li2018domain; carlucci2019; li2018domain; wang2019learning; akuzawa2020adversarial; zhu2022localized; gokhale2023improving; dayal2023madg; cheng2023adversarial; chen2024domain), and (ii) Approaches that construct an ensemble of domain-specific models, and then fuse their representations to form a unified, domain-agnostic representation (ding2017deep; mancini2018best; zhou2021domain; muhammad2024domain).
Additional strategies for domain generalization include contrastive learning approaches (kim2021selfreg), methods based on distribution alignment via metrics (muandet2013domain; sun2016deep), and techniques utilizing custom network architectures, for instance by incorporating domain-specific adapters between shared layers (rebuffi2017learning; rebuffi2018efficient; li2019efficient; omi2022model).